OpenStudy (anonymous):

simplify the expression and state any restrictions on the variables: 6m+9/6 i will give medals!

OpenStudy (mertsj):

\[\frac{6m+9}{6}=\frac{3(2m+3)}{6}=\frac{2m+3}{2}\]

OpenStudy (mertsj):

No restrictions on the variable